Sue paints a room at the rate of 1/10 room per hour, and Kelly paints the same room at the rate of 1/8 room per hour. How long will it take them to paint the room together?
Answer by checkley71(8403)

IT TAKES SUE 10 HOURS TO PAINT THE ROOM. 1/(1/10)=1*10/1=10
IT TAKES KELLY 8 HOURS TO PAINT THE ROOM. 1/(1/8)=1*8/1=8
10*8/(10+8)80/18=4.44 HOURS TO PAINT THE ROOM TOGETHER.